OpenStudy (anonymous):

Simplify. square root 8/^5 square root 8 A. 8^-7/10 B. 8^-1/4 C. 8^3/10 D. 8^7/10

OpenStudy (abdullahm):

Could you write that using the equation button or the draw tool? Thanks.

OpenStudy (anonymous):

\[\frac{ \sqrt{8} }{\sqrt[5]{8} ? }\] sorry the question mark is not supposed to be there lol

OpenStudy (abdullahm):

Hint: \(\sf\LARGE\sqrt[a]{b}=b^{\frac{1}{a}}\) Hint: \(\sf\LARGE\sqrt[]{c}=c^{\frac{1}{2}}\) Hint: \(\sf\Huge\frac{\sqrt[a]{b}}{\sqrt[c]{b}}=b^\left({\frac{1}{a}-\frac{1}{c}}\right)\)

OpenStudy (anonymous):

still dont get it :/

OpenStudy (abdullahm):

for those letters you see, just plug in numbers in their respective places then :P

OpenStudy (anonymous):

im sorry i still dont get it lol

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@bpstation

OpenStudy (anonymous):

im not good at square roots but khansacademy.com is

OpenStudy (anonymous):

ok thnx :)

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@vera_ewing

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@harryhart

OpenStudy (anonymous):

sorry i suck at math too

OpenStudy (anonymous):

lol its ok :)

OpenStudy (anonymous):

heres a calc: http://www.math.com/students/calculators/source/square-root.htm

OpenStudy (anonymous):

thnx :)